OpenStudy (anonymous):

Solve by completing the square. x^2 -20x=10 x=

OpenStudy (anonymous):

divide coeefficient of x by 2 and subtract 10^2 from both sides: (x - 10 )^2 -100 = 10-100 (x-10)^2 = 10 take sqrt of both sides and the rest is easy

OpenStudy (anonymous):

x^2 - 20x + 100 = 110 (x - 10)^2 = 110 x-10 = \[\sqrt{110}\] x=10\[\pm\]\[\sqrt{110}\]